OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 2. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (anonymous):

Do you know what slope-intercept form is?

OpenStudy (anonymous):

the slope-intercept form is y=mx+b where m is the gradient/slope and b is the y-intercept.

OpenStudy (anonymous):

Right. So the first step is to get the equation we have looking like that.

OpenStudy (anonymous):

ok but how do I do that

OpenStudy (anonymous):

Solve \(2x + 3y = 1200\) for y.

OpenStudy (anonymous):

400

OpenStudy (anonymous):

Er.. it has to be in the form of an equation. y = something.

OpenStudy (anonymous):

oh y=400

OpenStudy (anonymous):

What'd you do with x?

OpenStudy (anonymous):

I multiplied it by 0

OpenStudy (anonymous):

Ok, well, unless you want to solve for where x = 0, you shouldn't do that. \[3y = 1200 - 2x\]\[y = 400 - \frac{2}{3}x\]

OpenStudy (anonymous):

ooooo so would that be answer or is there more steps

OpenStudy (anonymous):

That's the equation in y=mx+b form. Now you know the slope and y intercept so you can graph it. The actual answer is to explain what we did, so... you'll have to do that. But it's basically the answer.

OpenStudy (anonymous):

okay thank you since you helped me with that can you tell me how to turn them into function notations

OpenStudy (anonymous):

What do you mean? You mean like f(x)?

OpenStudy (anonymous):

I don't know all it says is 3.Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences.

OpenStudy (anonymous):

Yeah. That's f(x). You just write the equation in terms of f(x).

OpenStudy (anonymous):

how do I do that

OpenStudy (anonymous):

Once you've solved for y (that is, it's in a form like \(y = mx + b\)) you just replace y with f(x).

OpenStudy (anonymous):

so fx= 400-2/3x then what because im supposed to graph the equation after I put it in a notation

OpenStudy (anonymous):

Ok. Well, you know the slope and the y-intercept. So you can graph it from that.

OpenStudy (anonymous):

wait whats the slope again

OpenStudy (anonymous):

\[y = mx + b\]m is the slope

OpenStudy (anonymous):

oh I thought m was the y intercept

OpenStudy (anonymous):

o wait nvm I got confused

OpenStudy (anonymous):

No, b is the y-intercept. (For some reason.)

OpenStudy (anonymous):

so since 2/3 is the slope how do I write that as a function

OpenStudy (anonymous):

do I do 2/3(400)

OpenStudy (anonymous):

I'm confused as to what you're asking.

OpenStudy (anonymous):

well because it wants me to write it as a notation then graph it so would the answer be 2/3(400). really I just need to know the notation so I can graph it

OpenStudy (anonymous):

The notation is \(f(x) = mx + b\) Where m is the slope and b is the y-intercept

OpenStudy (anonymous):

oooo okay I understand

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

just one last question and then im all done, the question is also apart of the function notation question. what does the graph of the function represent?

OpenStudy (anonymous):

I'm actually not sure what they're asking there. Profit of smart phone cases vs profit of tablet cases, I guess.

OpenStudy (anonymous):

Wait so the slope is 400 and the y intercept is 2/3? I'm confused here.